GIVE THE EQUATION FOR THE CIRCLE WITH CENTER C(3,-2) AND RADIUS 4.
Answer by uma(370)

The equation of the circle with centre C(3,-2) and radius 4 is given by:
(x-3)^2 + (y-(-2))^2 = 4^2
==> (x-3)^2 + (y+2)^2 = 16
On expanding we get,
x^2 - 6x + 9 + y^2 + 4y + 4 = 16
==> x^2 + y^2 - 6x + 4y + 13 = 16
==> x^2 + y^2 - 6x + 4y + 13 - 16 = 0 [adding - 16 to both the sides]
==> x^2 + y^2 - 6x + 4y - 3 = 0

Good Luck!!!